### v1

# Ivan Gatica and Alfonso Bustamante

Abstract
In this paper we present a brief study of the $\sigma$-set-$\sigma$-antiset duality that occurs in $\sigma$-set theory and we also present the development of the integer space $3^{A}=\left\langle 2^{A},2^{A^{-}}\right\rangle$ for the cardinals $|A|=2,3$ together with its algebraic properties. We also develop in this article a presentation of some of the properties of fusion of $\sigma$-sets and finally we present the development and definition of a type of equations of one $\sigma$-set variable.
1 $\sigma$-Sets and $\sigma$-Antisets
As we have seen in [6], an $\sigma$-antiset is defined as follows:
Definition 1.1.
Let $A$ be a $\sigma$-set, then $B$ is said to be the $\sigma$-antiset of $A$ if and only if $A\oplus B=\emptyset$, where $\oplus$ is the fusion of $\sigma$-sets.
We must observe that given the definition of the fusion operator $\oplus$ in [6] it is clear that it is commutative and therefore if $B$ is an $\sigma$-antiset of $A$, then it will be necessary that $A$ is also the $\sigma$-antiset of $B$. On the other hand, following Blizard notation, [4] p. 347, we will denote $B$ the $\sigma$-antiset of $A$ as $B=A^{-}$, in this way we will have that $A=(A^{-})^{-}$.
Continuing with the development of the $\sigma$-sets we have constructed three primary $\sigma$-sets, which are:
Natural Numbers =$\mathds{N}=\{1,2,3,4,5,6,7,8,9,10,\ldots\}$
00-Natural Numbers $\mathds{N}^{0}=\{1_{0},2_{0},3_{0},4_{0},5_{0},6_{0},7_{0},8_{0},9_{0},10_{0},\ldots\}$
Antinatural Numbers =$\mathds{N}^{-}=\{1^{\ast},2^{\ast},3^{\ast},4^{\ast},5^{\ast},6^{\ast},7^{\ast},8^{\ast},9^{\ast},10^{\ast},\ldots\}$
where $1=\{\alpha\}$, $1_{0}=\{\emptyset\}$ and $1^{\ast}=\{\omega\}$, we must clarify that we have changed the letter $\beta$ for the letter $\omega$ for symmetry reasons, we must also remember that:
$\ldots\in\alpha_{-2}\in\alpha_{-1}\in\alpha\in\alpha_{1}\in\alpha_{2}\ldots$
y
$\ldots\in\omega_{-2}\in\omega_{-1}\in\omega\in\omega_{1}\in\omega_{2}\in\ldots$
where both chains have the linear $\in$-root property and are totally different, i.e. they do not have a link-intersection.
On the other hand, we must remember the definition of the space generated by two $\sigma$-sets $A$ and $B$ which is:
Definition 1.2.
Let $A$ and $B$ be $\sigma$-sets. The Generated space by $A$ and $B$ is given by
$\left\langle 2^{A},2^{B}\right\rangle=\{x\oplus y:x\in 2^{A}\wedge y\in 2^{B}\},$
where $\oplus$ is the fusion operator.
Let us recall a few things about the fusion operator $\oplus$. To begin this brief analysis, we must observe that given $x$, $y$ two $\sigma$-sets, if $\{x\}\cup\{y\}=\emptyset$ then it will be said that $y$ is the antielement of $x$ and $x$ the antielement of $y$, where the union of pairs $\cup$ axiomatized within the theory of $\sigma$-sets is used, in particular in the completion axioms A and B, which we will call annihilation axioms from now on.
Notation 1.3.
Let $x$ be an element of some $\sigma$-set, then we will denote as $x^{\ast}$ the antielement of $x$, in the case that it exists.
Now we move on to define the new operations with $\sigma$-sets which will help us define the fusion of $\sigma$-sets.
Definition 1.4.
Let $A$ and $B$ be $\sigma$-sets, then we define the $\ast$-intersection of $A$ with $B$ by
$A\widehat{\cap}B=\{x\in A:x^{\ast}\in B\}.$
Example 1.5.
Let $A=\{1,2,3^{\ast},4\}$ and $B=\{2,3,4^{\ast}\}$, then we have that:
$A\widehat{\cap}B=\{3^{\ast},4\}$
and
$B\widehat{\cap}A=\{3,4^{\ast}\},$
it is clear that the $\ast$-intersection operator is not commutative.
Theorem 1.6.
Let $A$ be $\sigma$-set, then $A\widehat{\cap}A=\emptyset$.
Proof.
Let $A$ be a $\sigma$-set, by definition we will have that
$A\widehat{\cap}A=\{x\in A:x^{\ast}\in A\}.$
Suppose now that $A\widehat{\cap}A\neq\emptyset$, then there exists an $x\in A$ such that $x^{\ast}\in A$, therefore we will have that $x,x^{\ast}\in A$, which is a contradiction with Theorem 3.39 (Exclusion of inverses) from [6], so if $A$ is a $\sigma$-set then
$A\widehat{\cap}A=\emptyset.$
Regarding Theorem 1.6, we can observe that given a $\sigma$-set $A$ the $\sigma$-set theory does not allow the coexistence of a $\sigma$-element $x$ and its $\sigma$-antielement in the same $\sigma$-set $A$, and this is because $A$ is a $\sigma$-set. However, since $\sigma$-set theory is a $\sigma$-class theory, one can find the $\sigma$-elements together with the $\sigma$-antielements coexisting without problems in what we call the proper $\sigma$-class, in this way one will have that $\{x,x^{\ast}\}$ is a proper $\sigma$-class and not a $\sigma$-set.
Theorem 1.7.
Let $A$ be $\sigma$-set, then $A\widehat{\cap}\emptyset=\emptyset$ and ^$\emptyset\widehat{\cap}A=\emptyset$.
Proof.
Let $A$ be a $\sigma$-set, by definition we will have that
$A\widehat{\cap}\emptyset=\{x\in A:x^{\ast}\in\emptyset\}.$
Now suppose that $A\widehat{\cap}\emptyset\neq\emptyset$, then there exists an $x\in A$ such that $x^{\ast}\in\emptyset$, which is a contradiction, hence $A\widehat{\cap}\emptyset=\emptyset$. On the other hand, ^$\emptyset\widehat{\cap}A\subseteq\emptyset$ thus we will have that ^$\emptyset\widehat{\cap}A=\emptyset$. ∎
On the other hand, we will define the $\ast$-difference between $\sigma$-sets, a fundamental operation to be able to define the fusion between $\sigma$-sets.
Definition 1.8.
Let $A$ and $B$ be $\sigma$-sets, then we define the $\ast$-difference between $A$ y $B$ by
$A\divideontimes B=A-(A\widehat{\cap}B),$
where $A-B=\{x\in A:x\notin B\}.$
Example 1.9.
Let $A=\{1,2,3^{\ast},4\}$ and $B=\{2,3,4^{\ast}\}$, thenwe have that:
$A\widehat{\cap}B=\{3^{\ast},4\},$
therefore
$A\divideontimes B=A-(A\widehat{\cap}B)=\{1,2,3^{\ast},4\}-\{3^{\ast},4\}=\{1,2\}$
$A\divideontimes B=\{1,2\}.$
We also have to
$B\widehat{\cap}A=\{3,4^{\ast}\}$
therefore
$B\divideontimes A=B-(B\widehat{\cap}A)=\{2,3,4^{\ast}\}-\{3,4^{\ast}\}=\{2\}$
$B\divideontimes A=\{2\}.$
Corollary 1.10.
Let $A$ be $\sigma$-set. Then $A\divideontimes A=A$.
Proof.
Let $A$ be a $\sigma$-set, then by Theorem 1.6 we will have that $A\widehat{\cap}A=\emptyset$ therefore
$A\divideontimes A=A-(A\widehat{\cap}A)=A-\emptyset=A.$
Corollary 1.11.
Let $A$ be $\sigma$-set. Then $A\divideontimes\emptyset=A$ and $\emptyset\divideontimes A=\emptyset$.
Proof.
Let $A$ be a $\sigma$-set, then by Theorem 1.7 we will have that $A\widehat{\cap}\emptyset=\emptyset\widehat{\cap}A=\emptyset$ therefore
$A\divideontimes\emptyset=A-(A\widehat{\cap}\emptyset)=A-\emptyset=A$
and
$\emptyset\divideontimes A=\emptyset-(\emptyset\widehat{\cap}A)=\emptyset-\emptyset=\emptyset.$
Now after defining the $\ast$-intersection and the $\ast$-difference we can define the fusion of $\sigma$-sets as follows:
Definition 1.12.
Let $A$ and $B$ be $\sigma$-sets, then we define the fusion of $A$ and $B$ by
$A\oplus B=\{x:x\in A\divideontimes B\vee x\in B\divideontimes A\}.$
It is clear that the fusion of $\sigma$-sets is commutative by definition. Now let us show an example
Example 1.13.
Let $A=\{1,2,3^{\ast},4\}$ y $B=\{2,3,4^{\ast}\}$, then we have that:
$A\oplus B=\{x:x\in A\divideontimes B\vee x\in B\divideontimes A\},$
$A\oplus B=\{x:x\in\{1,2\}\vee x\in\{2\}\},$
$A\oplus B=\{1,2\},$
therefore we have that
$\{1,2,3^{\ast},4\}\oplus\{2,3,4^{\ast}\}=\{2,3,4^{\ast}\}\oplus\{1,2,3^{\ast},4\}=\{1,2\}.$
Corollary 1.14.
Let $A$ be a $\sigma$-set, then $A\oplus A=A$.
Proof.
Let $A$ be a $\sigma$-set, by definition we have that,
$A\oplus A=\{x:x\in A\divideontimes A\vee x\in A\divideontimes A\}.$
Now by corollary 1.10, we have that
$A\oplus A=\{x:x\in A\vee x\in A\},$
$A\oplus A=\{x:x\in A\},$
therefore it is clear that $A\subset A\oplus A$ and that $A\oplus A\subset A$, therefore $A\oplus A=A$.
Corollary 1.15.
Let $A$ be $\sigma$-set, then $A\oplus\emptyset=\emptyset\oplus A=A$.
Proof.
First we will show that $A\oplus\emptyset=A$. By definition we will have that,
$A\oplus\emptyset=\{x:x\in A\divideontimes\emptyset\vee x\in\emptyset\divideontimes A\}.$
Now by the corollary 1.11, we will have that
$A\oplus\emptyset=\{x:x\in A\vee x\in\emptyset\},$
$A\oplus\emptyset=\{x:x\in A\},$
from this it is clear that $A\subset A\oplus\emptyset$ and that $A\oplus\emptyset\subset A$, in this way $A\oplus\emptyset=A$.
Second we will show that $\emptyset\oplus A=A$. By definition we will have that,
$\emptyset\oplus A=\{x:x\in\emptyset\divideontimes A\vee x\in A\divideontimes\emptyset\}.$
Now by the corollary 1.11, we will have that
$\emptyset\oplus A=\{x:x\in\emptyset\vee x\in A\},$
$A\oplus\emptyset=\{x:x\in A\},$
from this it is clear that $A\subset\emptyset\oplus A$ and that $\emptyset\oplus A\subset A$, in this way $\emptyset\oplus A=A$. ∎
Theorem 1.16.
Let $X$ be a $\sigma$-set, then for all $A,B\in 2^{X}$, we have that:
$A\oplus B=A\cup B,$
where $A\cup B=\{x:x\in A\vee x\in B\}.$
Proof.
Let $X$ be a $\sigma$-set and $A,B\in 2^{X}$. Then, by theorem 3.39 of [6] we have that
$A\widehat{\cap}B=B\widehat{\cap}A=\emptyset,$
in this way
$A\divideontimes B=A\wedge B\divideontimes A=B.$
Finally $A\oplus B=\{x:x\in A\vee x\in B\}=A\cup B.$
Corollary 1.17.
Let $X$ be a $\sigma$-set, then for all $A\in 2^{X}$, we have that:
$A\oplus X=X.$
Proof.
Let $X$ be a $\sigma$-set and $A\in 2^{X}$. Then by theorem 1.16 we have that
$A\oplus X=A\cup X.$
Now as $A\subset X$, then $A\cup X=X$, therefore
$A\oplus X=X.$
As we said before the fusion of $\sigma$-sets $\oplus$ is commutative by definition but as we demonstrated in [6] this operation is not associative.
Example 1.18.
Let $A=\{1^{\ast},2^{\ast}\}$, $B=\{1,2\}$ y $C=\{1\}$, then
$(A\oplus B)\oplus C=\emptyset\oplus C=C$
and
$A\oplus(B\oplus C)=A\oplus B=\emptyset,$
therefore we have that
$(A\oplus B)\oplus C\neq A\oplus(B\oplus C).$
2 Generated space
As we have already indicated in the definition 1.2 we will have that the space generated by two $\sigma$-sets $A$ and $B$ is:
$\left\langle 2^{A},2^{B}\right\rangle=\{x\oplus y:x\in 2^{A}\wedge y\in 2^{B}\}.$
Now taking into account the duality $\sigma$-set, $\sigma$-antiset we could consider the following example.
Example 2.1.
We consider the $\sigma$-set $A=\{1,2,3\}$ and its $\sigma$-antiset $A^{-}=\{1^{\ast},2^{\ast},3^{\ast}\}$ then we obtain the integer space $3^{A}$ where,
$3^{A}=\left\langle 2^{A},2^{A^{-}}\right\rangle.$
Is important to observe that
$2^{A}=\{\emptyset,\{1\},\{2\},\{3\},\{1,2\},\{1,2\},\{2,3\},A\}$
and
$2^{A^{-}}=\{\emptyset^{-},\{1^{\ast}\},\{2^{\ast}\},\{3^{\ast}\},\{1^{\ast},2^{\ast}\},\{1^{\ast},2^{\ast}\},\{2^{\ast},3^{\ast}\},A^{-}\}.$
Also is important to observe that $\emptyset=\emptyset^{-}$, which is very important for the construction of $3^{A}$.
Now considering the definition of generated space,
$3^{A}=\left\langle 2^{A},2^{A^{-}}\right\rangle=\{X\oplus Y:X\in 2^{A}\wedge Y\in 2^{A^{-}}\},$
where the operator $\oplus$ is the fusion of $\sigma$-sets, we will obtain the following matrix:
$\oplus$ $\emptyset$ $\{1\}$$\{2\}$$\{3\}$$\{1,2\}$$\{1,3\}$$\{2,3\}$ $A$
$\emptyset^{-}$ $\emptyset^{0}_{0}$$\{1\}$$\{2\}$$\{3\}$$\{1,2\}$$\{1,3\}$$\{2,3\}$ $A$
$\{1^{\ast}\}$$\{1^{\ast}\}$ $\emptyset^{1}_{1}$$\{1^{\ast},2\}$$\{1^{\ast},3\}$$\{2\}$$\{3\}$ $\{1^{\ast},2,3\}$$\{2,3\}$
$\{2^{\ast}\}$$\{2^{\ast}\}$$\{1,2^{\ast}\}$ $\emptyset^{2}_{1}$$\{2^{\ast},3\}$$\{1\}$ $\{1,2^{\ast},3\}$$\{3\}$$\{1,3\}$
$\{3^{\ast}\}$$\{3^{\ast}\}$$\{1,3^{\ast}\}$$\{2,3^{\ast}\}$ $\emptyset^{3}_{1}$$\{1,2,3^{\ast}\}$$\{1\}$$\{2\}$$\{1,2\}$
$\{1^{\ast},2^{\ast}\}$$\{1^{\ast},2^{\ast}\}$$\{2^{\ast}\}$$\{1^{\ast}\}$$\{1^{\ast},2^{\ast},3\}$ $\emptyset^{4}_{2}$$\{2^{\ast},3\}$$\{1^{\ast},3\}$$\{3\}$
$\{1^{\ast},3^{\ast}\}$$\{1^{\ast},3^{\ast}\}$$\{3^{\ast}\}$$\{1^{\ast},2,3^{\ast}\}$$\{1^{\ast}\}$$\{2,3^{\ast}\}$ $\emptyset^{5}_{2}$$\{1^{\ast},3\}$$\{2\}$
$\{2^{\ast},3^{\ast}\}$$\{2^{\ast},3^{\ast}\}$$\{1,2^{\ast},3^{\ast}\}$$\{3^{\ast}\}$$\{2^{\ast}\}$$\{1,3^{\ast}\}$$\{1,2^{\ast}\}$ $\emptyset^{6}_{2}$$\{1\}$
$A^{-}$ $A^{-}$$\{2^{\ast},3^{\ast}\}$$\{1^{\ast},3^{\ast}\}$$\{1^{\ast},2^{\ast}\}$$\{3^{\ast}\}$$\{2^{\ast}\}$$\{1^{\ast}\}$ $\emptyset^{7}_{3}$
Table 1: Integer Space.
It is important to note that from the perspective of $\sigma$-sets we have that $\emptyset=\emptyset^{-}=\emptyset^{i}_{j}$ with $i\in\{0,1,2,3,4,5,6,7\}$ and $j\in\{0,1,2,3\}$, where the difference of the $\sigma$-emptysets $\emptyset^{i}_{j}$ is given by annihilation, which come from the equation $A\oplus A^{-}=\emptyset$.
From the matrix representation of the integer space $3^{A}$, we can present another representation of the same integer space. This representation of the integer space $3^{A}$ is a graphical representation which we show in figure 1.
Figure 1: Integer Space $3^{A}$.
Finally as a theoretical result we have a cardinal theorem:
Theorem 2.2.
Let $A=\{1,2,3\}$, then $\left|3^{A}\right|=\left|\left\langle 2^{A},2^{A^{-}}\right\rangle\right|=3^{3}=27$.
Proof.
Let $A=\{1,2,3\}$, the proof is the same fusion matrix for this $\sigma$-set. ∎
We should also note that we have obtained other cardinal results for the integer space $3^{A}$ with $\left|A\right|\in\{0,1,2,3,4,5\}$. The cardinal results are the following:
$\sigma$-Set $\sigma$-Antiset Generated Cardinal
$A=\emptyset$ $A^{-}=\emptyset^{-}$ $\left\langle 2^{A},2^{A^{-}}\right\rangle$ $13^{0}=$ 1
$A=\{1\}$ $A^{-}=\{1^{\ast}\}$ $\left\langle 2^{A},2^{A^{-}}\right\rangle$ $33^{1}=$ 3
$A=\{1,2\}$ $A^{-}=\{1^{\ast},2^{\ast}\}$ $\left\langle 2^{A},2^{A^{-}}\right\rangle$ $93^{2}=$ 9
$A=\{1,2,3\}$ $A^{-}=\{1^{\ast},2^{\ast},3^{\ast}\}$ $\left\langle 2^{A},2^{A^{-}}\right\rangle$ $273^{3}=$ 27
$A=\{1,2,3,4\}$ $A^{-}=\{1^{\ast},2^{\ast},3^{\ast},4^{\ast}\}$ $\left\langle 2^{A},2^{A^{-}}\right\rangle$ $813^{4}=$ 81
$A=\{1,2,3,4,5\}$ $A^{-}=\{1^{\ast},2^{\ast},3^{\ast},4^{\ast},5^{\ast}\}$ $\left\langle 2^{A},2^{A^{-}}\right\rangle$ $2433^{5}=$ 243
From these calculations made with the fusion matrix we can obtain the following conjecture.
Conjecture 2.3.
Let $A$ be $\sigma$-set such that $|A|=n$, then $\left|3^{A}\right|=\left|\left\langle 2^{A},2^{A^{-}}\right\rangle\right|=3^{n}$.
On the other hand, as we have already said, we are going to change the notation of $1_{\Theta}$ to $1_{0}$, in this way we will have the set 00-natural numbers defined as follows:
$1_{0}=\{\emptyset\}$
$2_{0}=\{\emptyset,1_{0}\}$
$3_{0}=\{\emptyset,1_{0},2_{0}\}$
$4_{0}=\{\emptyset,1_{0},2_{0},3_{0}\}$
and so on, forming the 00-natural numbers
$\mathds{N}^{0}=\{1_{0},2_{0},3_{0},4_{0},5_{0},6_{0},7_{0},8_{0},9_{0},10_{0},\ldots\},$
where one of the important properties of this $\sigma$-set is that it does not annihilate with the natural numbers $\mathds{N}$ nor with the antinatural numbers $\mathds{N}^{-}$, in this way we can consider the following example for the generated space.
Example 2.4.
We consider the $\sigma$-sets $A=\{1_{0},2_{0}\}$ and $B=\{1,2\}$,therefore the space generated by $A\oplus B$ and $A\oplus B^{-}$ will be:
$\left\langle 2^{A\oplus B},2^{A\oplus B^{-}}\right\rangle=\{x\oplus y:x\in 2^{A\oplus B}\wedge y\in 2^{A\oplus B^{-}}\}$
$\left\langle 2^{A\oplus B},2^{A\oplus B^{-}}\right\rangle=\{\emptyset,\{1_{0}\},\{1\},\{1^{\ast}\},\{2_{0}\},\{2\},\{2^{\ast}\},\{1_{0},2_{0}\},\{1_{0},1\},\{1_{0},1^{\ast}\},\{1_{0},2\},\{1_{0},2^{\ast}\},\\ \{2_{0},1\},\{2_{0},1^{\ast}\},\{2_{0},2\},\{2_{0},2^{\ast}\},\{1,2\},\{1,2^{\ast}\},\{1^{\ast},2\},\{1^{\ast},2^{\ast}\},\{1_{0},1,2\},\{1_{0},1,2^{\ast}\},\{1_{0},1^{\ast},2\},\\ \{1_{0},1^{\ast},2^{\ast}\},\{2_{0},1,2\},\{2_{0},1,2^{\ast}\},\{2_{0},1^{\ast},2\},\{2_{0},1^{\ast},2^{\ast}\},\{1_{0},2_{0},1\},\{1_{0},2_{0},1^{\ast}\},\{1_{0},2_{0},2\},\{1_{0},2_{0},2^{\ast}\},\\ \{1_{0},2_{0},1,2\},\{1_{0},2_{0},1,2^{\ast}\},\{1_{0},2_{0},1^{\ast},2\},\{1_{0},2_{0},1^{\ast},2^{\ast}\}\}$
In this case the generated space becomes a meta-space generated by $A=\{1_{0},2_{0}\}$ and $B=\{1,2\}$ which can be ordered graphically as shown in figure 2.
Figure 2: Meta-space $\left\langle 2^{A\oplus B},2^{A\oplus B^{-}}\right\rangle$.
Now if we count the number of elements that the meta-space generated by $A=\{1_{0},2_{0}\}$ and $B=\{1,2\}$ has, we will have that they are 36, where the prime decomposition of this number is $36=2^{2}\cdot 3^{2}$ which is equivalent to the following multiplication of cardinals $36=2^{|A|}\cdot 3^{|B|}$, from here we can obtain the following conjecture:
Conjecture 2.5.
For all $A\in 2^{\mathds{N}^{0}}$ and $B\in 2^{\mathds{N}}$, then $\left|\left\langle 2^{A\oplus B},2^{A\oplus B^{-}}\right\rangle\right|=2^{|A|}\cdot 3^{|B|}.$
Example 2.6.
We consider $A=\{1_{0}\}$ and $B=\{1,2\}$, then we obtain that
$\left\langle 2^{A\oplus B},2^{A\oplus B^{-}}\right\rangle=\{\emptyset,\{1_{0}\},\{1\},\{1^{\ast}\},\{2\},\{2^{\ast}\},\{1_{0},1\},\{1_{0},2\},\{1_{0},1^{\ast}\},\{1_{0},2^{\ast}\},\{1,2\},\{1,2^{\ast}\},\\ \{1^{\ast},2\},\{1^{\ast},2^{\ast}\},\{1_{0},1,2\},\{1_{0},1,2^{\ast}\},\{1_{0},1^{\ast},2\},\{1_{0},1^{\ast},2^{\ast}\}\}$
Thus, we have that $|A|=1$ and $|B|=2$ and $\left|\left\langle 2^{A\oplus B},2^{A\oplus B^{-}}\right\rangle\right|=2^{|A|}\cdot 3^{|B|}=2^{1}\cdot 3^{2}=18.$
Example 2.7.
We consider $A=\emptyset$ and $B=\{1,2\}$, then we obtain that
$3^{B}=\{\emptyset,\{1\},\{1^{\ast}\},\{2\},\{2^{\ast}\},\{1,2\},\{1,2^{\ast}\},\{1^{\ast},2\},\{1^{\ast},2^{\ast}\}\}$
Thus, we have that $|A|=0$ and $|B|=2$ and $\left|\left\langle 2^{A\oplus B},2^{A\oplus B^{-}}\right\rangle\right|=2^{|A|}\cdot 3^{|B|}=2^{0}\cdot 3^{2}=9.$
3 Algebraic structure of integer space $3^{A}$
With respect to the algebraic structure of the Integer Space $3^{A}$ for all $A\in 2^{\mathds{N}}$ we think that these structures are related with structures called NAFIL (non-associative finite invertible loops)
Theorem 3.1.
Let $A=\{1,2\}$, then $(3^{A},\oplus)$ satisfies the following conditions:
  1. 1.
    $(\forall X,Y\in 3^{A})(X\oplus Y\in 3^{A}),$
  2. 2.
    $(\exists!\emptyset\in 3^{A})(\forall X\in 3^{A})(X\oplus\emptyset=\emptyset\oplus X=X),$
  3. 3.
    $(\forall X\in 3^{A})(\exists!X^{-}\in 3^{A})(X\oplus X^{-}=X^{-}\oplus X=\emptyset),$
  4. 4.
    $(\forall X,Y\in 3^{A})(X\oplus Y=Y\oplus X).$
Proof.
Let $A=\{1,2\}$, then we quote the fusion matrix represented in table 2 for $3^{\{1,2\}}$.
From here it is clearly seen that the conditions $(1)$, $(2)$ and $(3)$ of theorem 3.1 are satisfied, where the condition $(4)$ is obvious by definition.
We must clarify that since $\sigma$-set $\emptyset=\emptyset^{-}$, and also $\emptyset=\emptyset^{0}_{0}=\emptyset^{1}_{1}=\emptyset^{2}_{1}=\emptyset^{3}_{2}$, from here we have the condition $(2)$ and the difference is in another dimension, the dimension of annihilation. Here we must clarify that the fusion operation $\oplus$ is not associative. Let $X=\{1^{\ast},2^{\ast}\}$, $Y=\{1,2\}$ and $Z=\{1\}$ then we will have that ($(\{1^{\ast},2^{\ast}\}\oplus\{1,2\})\oplus\{1\}=\emptyset\oplus\{1\}=\{1\}$
on the other hand
$\{1^{\ast},2^{\ast}\}\oplus(\{1,2\}\oplus\{1\})=\{1^{\ast},2^{\ast}\}\oplus\{1,2\}=\emptyset$
therefore we have that
$(X\oplus Y)\oplus Z\neq X\oplus(Y\oplus Z),$
which shows that the structure $(3^{A},\oplus)$, is non-associative.
$\oplus$ $\emptyset$ $\{1\}$$\{2\}$$\{1,2\}$
$\emptyset^{-}$ $\emptyset^{0}_{0}$$\{1\}$$\{2\}$$\{1,2\}$
$\{1^{\ast}\}$$\{1^{\ast}\}$ $\emptyset^{1}_{1}$$\{1^{\ast},2\}$$\{2\}$
$\{2^{\ast}\}$$\{2^{\ast}\}$$\{1,2^{\ast}\}$ $\emptyset^{2}_{1}$$\{1\}$
$\{1^{\ast},2^{\ast}\}$$\{1^{\ast},2^{\ast}\}$$\{2^{\ast}\}$$\{1^{\ast}\}$ $\emptyset^{3}_{2}$
Table 2: Integer Space $3^{\{1,2\}}$.
We now present a new conjecture.
Conjecture 3.2.
Let $A\in 2^{\mathds{N}}$, then $(3^{A},\oplus)$ satisfies the following conditions:
  1. 1.
    $(\forall X,Y\in 3^{A})(X\oplus Y\in 3^{A}),$
  2. 2.
    $(\exists!\emptyset\in 3^{A})(\forall X\in 3^{A})(X\oplus\emptyset=\emptyset\oplus X=X),$
  3. 3.
    $(\forall X\in 3^{A})(\exists!X^{-}\in 3^{A})(X\oplus X^{-}=X^{-}\oplus X=\emptyset),$
  4. 4.
    $(\forall X,Y\in 3^{A})(X\oplus Y=Y\oplus X).$
4 $\sigma$-Sets Equations
Continuing with the analysis of the $\sigma$-sets, we now have the development of the equations of $\sigma$-sets of a $\sigma$-set variable, equations that play a very important role when solving a $\sigma$-set equation, now let’s define and go deeper into the $\sigma$-sets variables.
We must remember that for every $\sigma$-set $A$ and $B$, the fusion of both is defined as:
$A\oplus B=\{x:x\in A\divideontimes B\vee x\in B\divideontimes A\}$
Definition 4.1.
Let $A$ be a $\sigma$-set, then $A$ is said to be an entire $\sigma$-set if there exists the $\sigma$-antiset $A^{-}$.
Example 4.2.
Let $A=\{1_{0},2_{0},3_{0}\}$, then this $\sigma$-set is not an integer, since $A^{-}$ does not exist, on the other hand the $\sigma$-set $A=\{1,2,3,4\}$, is an integer $\sigma$-set since $A^{-}=\{1^{\ast},2^{\ast},3^{\ast},4^{\ast}\}$ exists which is the $\sigma$-antiset of $A$.
It is clear that if a $\sigma$-set $A$ is integer, then by definition there exists the integer space $3^{A}$. We should also note that if $A$ is an integer $\sigma$-set, then $\left[A\cup A^{-}\right]$ is a proper $\sigma$-class, for example, consider $A=\{1,2\}$, then $\left[A\cup A^{-}\right]=\left[1,2,1^{\ast},2^{\ast}\right]$, is a proper $\sigma$-class.
Definition 4.3.
Let $A$ be an integer $\sigma$-set such that $|A|=n$, then $X$ is said to be a $\sigma$-set variable of $3^{A}$, if and only if
$X=\{x_{1},x_{2},x_{3},\ldots,x_{m}\},$
where $m\leq n$ and $x_{i}$ a variable of the proper class $\left[A\cup A^{-}\right]$.
Example 4.4.
Let $A=\{1,2,3\}$ be a $\sigma$-set, it is clear that $A$ is an entire $\sigma$-set since there exists $A^{-}=\{1^{\ast},2^{\ast},3^{\ast}\}$ and therefore $3^{A}$, in this way we will have that
$X=\emptyset,$
$X=\{x\},$
$X=\{x_{1},x_{2}\},$
$X=\{x_{1},x_{2},x_{3}\},$
are $\sigma$-sets variables of $3^{A}$, where $x,x_{1},x_{2},x_{3}\in\left[1,2,3,1^{\ast},2^{\ast},3^{\ast}\right]$.
Lemma 4.5.
Let $A$ be an integer $\sigma$-set and $X$ a $\sigma$-set variable of $3^{A}$, then $A\oplus X=A\cup X$, with $A\subset A\cup X$ and $X\subset A\cup X$.
Proof.
Let $A$ be an integer $\sigma$-set and $X$ a $\sigma$-set variable of $3^{A}$, then
$A\oplus X=\{x:x\in A\divideontimes X\vee x\in X\divideontimes A\}$
Now we have that
$A\divideontimes X=A$
and
$X\divideontimes A=X$
since $X$ is a $\sigma$-set variable, therefore we will have that
$A\oplus X=\{x:x\in A\vee x\in X\}=A\cup X.$
We can also observe that $A\cap X=\emptyset$ since $X$ is a $\sigma$-set variable, therefore $A\subset A\cup X$ and $X\subset A\cup X$. ∎
Example 4.6.
Let $A=\{1,2,3\}$, and $X$ be a $\sigma$-set variable of $3^{A}$, that is,
$X=\emptyset$
$X=\{x\},$
$X=\{x_{1},x_{2}\},$
$X=\{x_{1},x_{2},x_{3}\},$
are $\sigma$-sets variable of $3^{A}$, where $x,x_{1},x_{2},x_{3}\in\left[1,2,3,1^{\ast},2^{\ast},3^{\ast}\right]$. then
$A\oplus X=\{1,2,3\},$
$A\oplus X=\{1,2,3,x\},$
$A\oplus X=\{1,2,3,x_{1},x_{2}\},$
$A\oplus X=\{1,2,3,x_{1},x_{2},x_{3}\}$
After the lemma 4.5 we proceed to analyze some equations of a $\sigma$-set variable and their solutions
Let $A$ be an integer $\sigma$-set, $X$ a $\sigma$-set variable and $M$ and $N$ two $\sigma$-sets of the integer space $3^{A}$, then an equation of a $\sigma$-set variable will be
$X\oplus M=N.$
Now if $M=N$, then the equation becomes
$X\oplus M=M,$
and by the corollary 1.17 we have that the solutions are all $X\in 2^{M}$, where we naturally count $X=\emptyset$, hence we have an equation of a $\sigma$-set variable with multiple solutions.
Now consider $M\neq N$, then the $\sigma$-set equation becomes:
$X\oplus M=N,$
We must remember that the structure in general is not associative, therefore we cannot freely use this property, so to find the solution to the equation we must develop a previous theorem. To develop this theorem we will assume that for every integer $\sigma$-set $A$ then the generated space $\left\langle 2^{A},2^{A^{-}}\right\rangle=3^{A}$, and also that $3^{A}$ satisfies the conjecture 3.2.
Theorem 4.7.
Let $A$ be an integer $\sigma$-set, $X$ be a $\sigma$-set variable of $3^{A}$ and $M\in 3^{A}$. Then
$(X\oplus M)\oplus M^{-}=X$
Proof.
Let $A$ be an entire $\sigma$-set, $X$ be a $\sigma$-set variable of $3^{A}$ and $M\in 3^{A}$, then by lemma 4.5 we have that $X\oplus M=X\cup M$, with $X\cap M=\emptyset$.
Therefore we have that
$\circledast\ (X\oplus M)\oplus M^{-}=\{a:a\in(X\oplus M)\divideontimes M^{-}\vee a\in M^{-}\divideontimes(X\oplus M)\}$
=$=\{a:a\in(X\cup M)\divideontimes M^{-}\vee a\in M^{-}\divideontimes(X\cup M)\}$
so
$(X\cup M)\divideontimes M^{-}=(X\cup M)-(X\cup M)\widehat{\cap}M^{-}=(X\cup M)-M=X,$
and
$M^{-}\divideontimes(X\cup M)=M^{-}-M^{-}\widehat{\cap}(X\cup M)=M^{-}-M^{-}=\emptyset.$
[…]
$(X\oplus M)\oplus M^{-}=\{a:a\in X\vee a\in\emptyset\}$
$(X\oplus M)\oplus M^{-}=\{a:a\in X\},$
$(X\oplus M)\oplus M^{-}=X.$
Now that theorem 4.7 has been proved, we can solve some $\sigma$-set equation for the integer $\sigma$-set $A=\{1,2\}$, since the generated space is effectively equal to $3^{A}$, that is, $\left\langle 2^{A},2^{A^{-}}\right\rangle=3^{A}$, and also $3^{A}$ is a non-associative abelian loop.
Let $A=\{1,2\}$ be an integer set and $M,N\in 3^{A}$, then the equation
$X\oplus M=N,$
has the following solution
$X\oplus M=N\setminus\oplus M^{-},$
$(X\oplus M)\oplus M^{-}=N\oplus M^{-},$
then by theorem 4.7 we will have that
$X=N\oplus M^{-}.$
Let us now show a concrete example for $A=\{1,2\}$.
Example 4.8.
Let $A=\{1,2\}$ be an integer $\sigma$-set, $M=\{1,2^{\ast}\}$ and $N=\{1\}$, then the equation of a $\sigma$-set variable
$X\oplus\{1,2^{\ast}\}=\{1\}$
has the following solution.
$X\oplus\{1,2^{\ast}\}=\{1\}\setminus\oplus\{1^{\ast},2\},$
$(X\oplus\{1,2^{\ast}\})\oplus\{1^{\ast},2\}=\{1\}\oplus\{1^{\ast},2\},$
$X=\{2\}.$
Here we can see that the equation has as solution the $\sigma$-set $S_{1}=\{2\}$, since
$\{2\}\oplus\{1,2^{\ast}\}=\{1\},$
but like the equation $X\oplus M=M$, this one does not have a unique solution since the $\sigma$-set $S_{2}=\{1,2\}$, is also a solution for the equation of a $\sigma$-set variable,
$\{1,2\}\oplus\{1,2^{\ast}\}=\{1\}.$
In this way we have two solutions for our equation of a $\sigma$-set variable which are:
$R=\{S_{1},S_{2}\}=\{\{2\},\{1,2\}\}.$
To conclude our study of $\sigma$-sets and $\sigma$-antisets we are going to present a conjecture regarding the possible solutions of a particular type of equations of a $\sigma$-set variable.
Conjecture 4.9.
Let $A$ be an integer $\sigma$-set, $X$ a $\sigma$-set variable of $3^{A}$, and $M,N\in 3^{A}$, then a solution of the equation
$X\oplus M=N,$
is $S=N\oplus M^{-}$.

### v2

@@
 In this paper we present a brief study of the $\sigma$-set-$\sigma$-antiset duality that occurs in $\sigma$-set theory and we also present the development of the integer space $3^{A}=\left\langle 2^{A},2^{A^{-}}\right\rangle$ for the cardinals $|A|=2,3$ together with its algebraic properties.
-We also develop in this article a presentation of some of the properties of fusion of $\sigma$-sets and finally we present the development and definition of a type of equations of one $\sigma$-set variable.
+In this article, we also develop a presentation of some of the properties of fusion of $\sigma$-sets and finally we present the development and definition of a type of equations of one $\sigma$-set variable.
 1 $\sigma$-Sets and $\sigma$-Antisets
@@
 $A\oplus X=X.$
-As we said before the fusion of $\sigma$-sets $\oplus$ is commutative by definition but as we demonstrated in [6] this operation is not associative.
+As we said before the fusion of $\sigma$-sets $\oplus$ is commutative by definition but as we demonstrated in [6, 3, 2] this operation is not associative.
 Example 1.18.
@@
 In this way we have two solutions for our equation of a $\sigma$-set variable which are:
-$R=\{S_{1},S_{2}\}=\{\{2\},\{1,2\}\}.$
-To conclude our study of $\sigma$-sets and $\sigma$-antisets we are going to present a conjecture regarding the possible solutions of a particular type of equations of a $\sigma$-set variable.
+$S=\{S_{1},S_{2}\}=\{\{2\},\{1,2\}\}.$
+Before we conclude our study of $\sigma$-sets and $\sigma$-antisets we are going to present a conjecture regarding the possible solutions of a particular type of equations of a $\sigma$-set variable.
 Conjecture 4.9.
@@
 is $S=N\oplus M^{-}$.
+Let us look into the following case, where $M=\{1^{*}\}$ and $N=\{1\}$, so for the equation
+$X\oplus\{1^{*}\}=\{1\},$
+there is no element $x$ such that $\{x,1^{*}\}=\{1\}$.
+This lead us into a more general case where $N\subseteq M^{-}$ or $N\supset M^{-}$, being the conjecture false, which take us into establishing the following condition in order for this type of equation being solvable.
+Definition 4.10.
+A $\sigma$-set equation $X\oplus M=N$ is said to be fusionable if $M\widehat{\cap}N=\emptyset$.
+With this, let us conclude with a bounded theorem to find some solutions of the $\sigma$-set equation.
+Theorem 4.11.
+Let $A$ be an integer $\sigma$-set, $X$ a $\sigma$-set variable of $3^{A}$, and $M,N\in 3^{A}$, then two possible solutions $S=\{S_{1},S_{2}\}$ of the fusionable equation
+$X\oplus M=N,$
+are $S_{1}=N\oplus R^{-}$ and $S_{2}=R^{-}$, where $R:=M\oplus N^{-}.$
+Proof.
+For the first solution $S_{1}$ we have that
+$S_{1}$ $=$ $N\oplus R^{-}$
+$=$ $N\oplus(M\oplus N^{-})^{-}$
+$=$ $N\oplus(N\oplus M^{-})$
+$=$ $N\oplus M^{-},$
+where $S_{2}=(M\oplus N^{-})^{-}=N\oplus M^{-}$ because of the result iteration seen above.
+Hence both results are actually a fusion solution for $X\oplus M=N,$ where $S_{2}=R^{-}$ is an exact solution and $S_{1}=N\oplus R^{-}$ is an intersected rest solution.
+Because of $M\widehat{\cap}N=\emptyset$ (Definition 4.10) as the equation $X\oplus M=N$ is fusionable, both $S_{1}\oplus M$ and $S_{2}\oplus M$ will be fusionable into another $\sigma$-set $N$.
+∎
+As we looked above, the solution space is reduced such that the solutions are indeed $N\oplus M^{-},$ being by consequence possible solutions for the fusionable equation $X\oplus M=N.$
+Example 4.12.
+Let $A=\{1,2,3,4,5,6\}$ be an integer $\sigma$-set, $M=\{1,2,3^{*},4^{*},5,6^{*}\}$ and $N=\{1,2\}$, then the equation of a $\sigma$-set variable
+$X\oplus\{1,2,3^{*},4^{*},5,6^{*}\}=\{1,2\},$
+which is fusionable because $M\widehat{\cap}N=\{1,2,3^{*},4^{*},5,6^{*}\}\widehat{\cap}\{1,2\}=\emptyset.$
+Now, by using Theorem 4.11, let us first obtain
+$R$ $=$ $(M\oplus N^{-})^{-}$
+$=$ ($(\{1,2,3^{*},4^{*},5,6^{*}\}\oplus\{1,2\}^{-})^{-}$
+$=$ ($(\{1,2,3^{*},4^{*},5,6^{*}\}\oplus\{1^{*},2^{*}\})^{-}$
+$=$ ($(\{3^{*},4^{*},5,6^{*}\})^{-}$
+$=$ $\{3,4,5^{*},6\},$
+so we get $S_{1}=N\oplus R^{-}=\{1,2,3,4,5^{*},6\}$ and $S_{2}=R^{-}=\{3,4,5^{*},6\}$, which can be easily proved that both solutions gives $S_{1}\oplus M=S_{2}\oplus M=N$ as a resulting $\sigma$-set.
+Hence $S=\{\{1,2,3,4,5^{*},6\},\{3,4,5^{*},6\}\}$ is a solution set for the fusionable equation $X\oplus M=N$.

### v3

@@
 We must observe that given the definition of the fusion operator $\oplus$ in [5] it is clear that it is commutative and therefore if $B$ is an $\sigma$-antiset of $A$, then it will be necessary that $A$ is also the $\sigma$-antiset of $B$.
-On the other hand, following Blizard notation, [4] p.
-347, we will denote $B$ the $\sigma$-antiset of $A$ as $B=A^{-}$, in this way we will have that $A=(A^{-})^{-}$.
+On the other hand, following the Blizard notation, [3] p.
+347, we will denote $B$ the $\sigma$-antiset of $A$ as $B=A^{-}$, in this way we will have $A=(A^{-})^{-}$.
 Continuing with the development of the $\sigma$-sets we have constructed three primary $\sigma$-sets, which are:
@@
 $\ldots\in\alpha_{-2}\in\alpha_{-1}\in\alpha\in\alpha_{1}\in\alpha_{2}\ldots$
-y
+and
 $\ldots\in\omega_{-2}\in\omega_{-1}\in\omega\in\omega_{1}\in\omega_{2}\in\ldots$
-where both chains have the linear $\in$-root property and are totally different, i.e. they do not have a link-intersection.
+where both $\epsilon$-chains have the linear $\epsilon$-root property and are totally different, i.e.
+they do not have a link-intersection.
+These definitions can be found in [5] Definition 3.13, 3.14 and 3.16.
 On the other hand, we must remember the definition of the space generated by two $\sigma$-sets $A$ and $B$ which is:
 Definition 1.2.
-Let $A$ and $B$ be $\sigma$-sets. The Generated space by $A$ and $B$ is given by
+Let $A$ and $B$ be two $\sigma$-sets. The Generated space by $A$ and $B$ is given by
 $\left\langle 2^{A},2^{B}\right\rangle=\{x\oplus y:x\in 2^{A}\wedge y\in 2^{B}\},$
@@
 Let us recall a few things about the fusion operator $\oplus$.
-To begin this brief analysis, we must observe that given $x$, $y$ two $\sigma$-sets, if $\{x\}\cup\{y\}=\emptyset$ then it will be said that $y$ is the antielement of $x$ and $x$ the antielement of $y$, where the union of pairs $\cup$ axiomatized within the theory of $\sigma$-sets is used, in particular in the completion axioms A and B, which we will call annihilation axioms from now on.
+In this brief analysis, we must observe that given $x$, $y$ two $\sigma$-sets, if $\{x\}\cup\{y\}=\emptyset$ then it will be said that $y$ is the antielement of $x$ and $x$ the antielement of $y$, where the union of pairs $\cup$ axiomatized within the theory of $\sigma$-sets is used, in particular in the completion axioms A and B, which we will call annihilation axioms from now on.
 Notation 1.3.
-Let $x$ be an element of some $\sigma$-set, then we will denote as $x^{\ast}$ the antielement of $x$, in the case that it exists.
-Now we move on to define the new operations with $\sigma$-sets which will help us define the fusion of $\sigma$-sets.
+Let $x$ be an element of some $\sigma$-set, then we will denote by $x^{\ast}$ the anti-element of $x$, if it exists.
+Now we move on to define the new operations with $\sigma$-sets which will help us define the fusion of $\sigma$-sets $\oplus$.
 Definition 1.4.
-Let $A$ and $B$ be $\sigma$-sets, then we define the $\ast$-intersection of $A$ with $B$ by
+Let $A$ and $B$ be two $\sigma$-sets, then we define the $\ast$-intersection of $A$ with $B$ by
 $A\widehat{\cap}B=\{x\in A:x^{\ast}\in B\}.$
 Example 1.5.
-Let $A=\{1,2,3^{\ast},4\}$ and $B=\{2,3,4^{\ast}\}$, then we have that:
+Let $A=\{1,2,3^{\ast},4\}$ and $B=\{2,3,4^{\ast}\}$ be two $\sigma$-sets, then we have that:
 $A\widehat{\cap}B=\{3^{\ast},4\}$
@@
 Theorem 1.6.
-Let $A$ be $\sigma$-set, then $A\widehat{\cap}A=\emptyset$.
+Let $A$ be a $\sigma$-set, then $A\widehat{\cap}A=\emptyset$.
 Proof.
@@
 $A\widehat{\cap}A=\emptyset.$
-Regarding Theorem 1.6, we can observe that given a $\sigma$-set $A$ the $\sigma$-set theory does not allow the coexistence of a $\sigma$-element $x$ and its $\sigma$-antielement in the same $\sigma$-set $A$, and this is because $A$ is a $\sigma$-set.
+Example 1.7.
+Let $A=\{1,2,3^{*},4\}$, then
+$A\widehat{\cap}A=\{1,2,3^{*},4\}\widehat{\cap}\{1,2,3^{*},4\},$
+$A\widehat{\cap}A=\{x\in\{1,2,3^{*},4\}:x^{*}\in\{1,2,3^{*},4\}\},$
+$A\widehat{\cap}A=\emptyset.$
+Regarding Theorem 1.6, we can observe that given a $\sigma$-set $A$, the $\sigma$-set theory does not allow the coexistence of a $\sigma$-element $x$ and its $\sigma$-antielement in the same $\sigma$-set $A$, and this is because $A$ is a $\sigma$-set.
 However, since $\sigma$-set theory is a $\sigma$-class theory, one can find the $\sigma$-elements together with the $\sigma$-antielements coexisting without problems in what we call the proper $\sigma$-class, in this way one will have that $\{x,x^{\ast}\}$ is a proper $\sigma$-class and not a $\sigma$-set.
 Theorem 1.8.
-Let $A$ be $\sigma$-set, then $A\widehat{\cap}\emptyset=\emptyset$ and ^$\emptyset\widehat{\cap}A=\emptyset$.
+Let $A$ be a $\sigma$-set, then $A\widehat{\cap}\emptyset=\emptyset$ and ^$\emptyset\widehat{\cap}A=\emptyset$.
 Proof.
@@
 Now suppose that $A\widehat{\cap}\emptyset\neq\emptyset$, then there exists an $x\in A$ such that $x^{\ast}\in\emptyset$, which is a contradiction, hence $A\widehat{\cap}\emptyset=\emptyset$.
-On the other hand, ^$\emptyset\widehat{\cap}A\subseteq\emptyset$ thus we will have that ^$\emptyset\widehat{\cap}A=\emptyset$.
+On the other hand, ^$\emptyset\widehat{\cap}A\subseteq\emptyset$ thus we will have to ^$\emptyset\widehat{\cap}A=\emptyset$.
 ∎
@@
 Definition 1.9.
-Let $A$ and $B$ be $\sigma$-sets, then we define the $\ast$-difference between $A$ y $B$ by
+Let $A$ and $B$ be two $\sigma$-sets, then we define the $\ast$-difference between $A$ y $B$ by
 $A\divideontimes B=A-(A\widehat{\cap}B),$
@@
 Corollary 1.11.
-Let $A$ be $\sigma$-set. Then $A\divideontimes A=A$.
+Let $A$ be a $\sigma$-set. Then $A\divideontimes A=A$.
 Proof.
@@
 Corollary 1.12.
-Let $A$ be $\sigma$-set. Then $A\divideontimes\emptyset=A$ and $\emptyset\divideontimes A=\emptyset$.
+Let $A$ be a $\sigma$-set. Then $A\divideontimes\emptyset=A$ and $\emptyset\divideontimes A=\emptyset$.
 Proof.
@@
 Definition 1.13.
-Let $A$ and $B$ be $\sigma$-sets, then we define the fusion of $A$ and $B$ by
+Let $A$ and $B$ be two $\sigma$-sets, then we define the fusion of $A$ and $B$ by
 $A\oplus B=\{x:x\in A\divideontimes B\vee x\in B\divideontimes A\}.$
-It is clear that the fusion of $\sigma$-sets is commutative by definition. Now let us show an example
+It is clear that the fusion of $\sigma$-sets is commutative by definition. Now, let us show an example
 Example 1.14.
@@
 Corollary 1.16.
-Let $A$ be $\sigma$-set, then $A\oplus\emptyset=\emptyset\oplus A=A$.
+Let $A$ be a $\sigma$-set, then $A\oplus\emptyset=\emptyset\oplus A=A$.
 Proof.
@@
 from this it is clear that $A\subset A\oplus\emptyset$ and that $A\oplus\emptyset\subset A$, in this way $A\oplus\emptyset=A$.
-Second we will show that $\emptyset\oplus A=A$. By definition we will have that,
+Second, we will show that $\emptyset\oplus A=A$. By definition we will have that,
 $\emptyset\oplus A=\{x:x\in\emptyset\divideontimes A\vee x\in A\divideontimes\emptyset\}.$
@@
 Finally $A\oplus B=\{x:x\in A\vee x\in B\}=A\cup B.$
+Example 1.18.
+Let $X=\{1,2,3\}$, $A=\{1,2\}$ and $B=\{2,3\}$, it is clear that $A,B\in 2^{X}$. Now we apply the fusion operator $\oplus$.
+$A\oplus B=\{x:x\in A\divideontimes B\vee x\in B\divideontimes A\},$
+$A\oplus B=\{x:x\in A\vee x\in B\},$
+$A\oplus B=A\cup B=\{1,2,3\}.$
 Corollary 1.19.
@@
 $A\oplus X=X.$
-As we said before the fusion of $\sigma$-sets $\oplus$ is commutative by definition but as we demonstrated in [6, 3, 2] this operation is not associative.
+Example 1.20.
+Let $X=\{1,2,3,4\}$ and $A=\{1,2,3\}$, it is clear that $A\in 2^{X}$. Now we apply the fusion operator $\oplus$.
+$A\oplus X=\{x:x\in A\divideontimes X\vee x\in X\divideontimes A\},$
+$A\oplus B=\{x:x\in A\vee x\in X\},$
+$A\oplus X=A\cup X=\{1,2,3,4\}=X.$
+As we said before, the fusion of $\sigma$-sets $\oplus$ is commutative by definition but as we demonstrated in [5, 2, 1] this operation is not associative.
 Example 1.21.
@@
 Table 1: Integer Space.
-It is important to note that from the perspective of $\sigma$-sets we have that $\emptyset=\emptyset^{-}=\emptyset^{i}_{j}$ with $i\in\{0,1,2,3,4,5,6,7\}$ and $j\in\{0,1,2,3\}$, where the difference of the $\sigma$-emptysets $\emptyset^{i}_{j}$ is given by annihilation, which come from the equation $A\oplus A^{-}=\emptyset$.
+It is important to note that from the perspective of $\sigma$-sets we have that $\emptyset=\emptyset^{-}=\emptyset^{i}_{j}$ with $i\in\{0,1,2,3,4,5,6,7\}$ and $j\in\{0,1,2,3\}$, where the difference of the $\sigma$-emptysets $\emptyset^{i}_{j}$ is given by annihilation, which comes from equation $A\oplus A^{-}=\emptyset$.
 From the matrix representation of the integer space $3^{A}$, we can present another representation of the same integer space.
@@
 Figure 1: Integer Space $3^{A}$.
-Finally as a theoretical result we have a cardinal theorem:
+Finally, as a theoretical result, we have a cardinal theorem:
 Theorem 2.2.
@@
 Let $A=\{1,2,3\}$, the proof is the same fusion matrix for this $\sigma$-set. ∎
-We should also note that we have obtained other cardinal results for the integer space $3^{A}$ with $\left|A\right|\in\{0,1,2,3,4,5\}$. The cardinal results are the following:
+We should also note that we have obtained other cardinal results for the integer space $3^{A}$ with $\left|A\right|\in\{0,1,2,3,4,5\}$. The cardinal results are as follows:
 $\sigma$-Set $\sigma$-Antiset Generated Cardinal
@@
 Conjecture 2.3.
-Let $A$ be $\sigma$-set such that $|A|=n$, then $\left|3^{A}\right|=\left|\left\langle 2^{A},2^{A^{-}}\right\rangle\right|=3^{n}$.
-On the other hand, as we have already said, we are going to change the notation of $1_{\Theta}$ to $1_{0}$, in this way we will have the set 00-natural numbers defined as follows:
+Let $A$ be a $\sigma$-set such that $|A|=n$, then $\left|3^{A}\right|=\left|\left\langle 2^{A},2^{A^{-}}\right\rangle\right|=3^{n}$.
+On the other hand, as we have already said, we are going to change the notation of $1_{\Theta}$ to $1_{0}$, in this way we will have the $\sigma$-set of 00-natural numbers defined as follows:
 $1_{0}=\{\emptyset\}$
@@
 $\left\langle 2^{A\oplus B},2^{A\oplus B^{-}}\right\rangle=\{\emptyset,\{1_{0}\},\{1\},\{1^{\ast}\},\{2_{0}\},\{2\},\{2^{\ast}\},\{1_{0},2_{0}\},\{1_{0},1\},\{1_{0},1^{\ast}\},\{1_{0},2\},\{1_{0},2^{\ast}\},\\ \{2_{0},1\},\{2_{0},1^{\ast}\},\{2_{0},2\},\{2_{0},2^{\ast}\},\{1,2\},\{1,2^{\ast}\},\{1^{\ast},2\},\{1^{\ast},2^{\ast}\},\{1_{0},1,2\},\{1_{0},1,2^{\ast}\},\{1_{0},1^{\ast},2\},\\ \{1_{0},1^{\ast},2^{\ast}\},\{2_{0},1,2\},\{2_{0},1,2^{\ast}\},\{2_{0},1^{\ast},2\},\{2_{0},1^{\ast},2^{\ast}\},\{1_{0},2_{0},1\},\{1_{0},2_{0},1^{\ast}\},\{1_{0},2_{0},2\},\{1_{0},2_{0},2^{\ast}\},\\ \{1_{0},2_{0},1,2\},\{1_{0},2_{0},1,2^{\ast}\},\{1_{0},2_{0},1^{\ast},2\},\{1_{0},2_{0},1^{\ast},2^{\ast}\}\}$
-In this case the generated space becomes a meta-space generated by $A=\{1_{0},2_{0}\}$ and $B=\{1,2\}$ which can be ordered graphically as shown in figure 2.
+In this case, the generated space becomes a meta-space generated by $A=\{1_{0},2_{0}\}$ and $B=\{1,2\}$ which can be ordered graphically as shown in figure 2.
 Figure 2: Meta-space $\left\langle 2^{A\oplus B},2^{A\oplus B^{-}}\right\rangle$.
-Now if we count the number of elements that the meta-space generated by $A=\{1_{0},2_{0}\}$ and $B=\{1,2\}$ has, we will have that they are 36, where the prime decomposition of this number is $36=2^{2}\cdot 3^{2}$ which is equivalent to the following multiplication of cardinals $36=2^{|A|}\cdot 3^{|B|}$, from here we can obtain the following conjecture:
+Now, if we count the number of elements that the meta-space generated by $A=\{1_{0},2_{0}\}$ and $B=\{1,2\}$ has, we will find that they are 36, where the prime decomposition of this number is $36=2^{2}\cdot 3^{2}$ which is equivalent to the following multiplication of cardinals $36=2^{|A|}\cdot 3^{|B|}$, from where we can obtain the following conjecture:
 Conjecture 2.5.
@@
 Let $A=\{1,2\}$, then we quote the fusion matrix represented in table 2 for $3^{\{1,2\}}$.
-From here it is clearly seen that the conditions $(1)$, $(2)$ and $(3)$ of theorem 3.1 are satisfied, where the condition $(4)$ is obvious by definition.
-We must clarify that since $\sigma$-set $\emptyset=\emptyset^{-}$, and also $\emptyset=\emptyset^{0}_{0}=\emptyset^{1}_{1}=\emptyset^{2}_{1}=\emptyset^{3}_{2}$, from here we have the condition $(2)$ and the difference is in another dimension, the dimension of annihilation.
+From here it is clearly seen that conditions $(1)$, $(2)$, and $(3)$ of theorem 3.1 are satisfied, where the condition $(4)$ is obvious by definition.
+We must clarify that since $\sigma$-set $\emptyset=\emptyset^{-}$, and also $\emptyset=\emptyset^{0}_{0}=\emptyset^{1}_{1}=\emptyset^{2}_{1}=\emptyset^{3}_{2}$, from here we have condition $(2)$ and the difference is in another dimension, the dimension of annihilation.
 Here we must clarify that the fusion operation $\oplus$ is not associative.
@@
 We should also note that if $A$ is an integer $\sigma$-set, then $\left[A\cup A^{-}\right]$ is a proper $\sigma$-class, for example, consider $A=\{1,2\}$, then $\left[A\cup A^{-}\right]=\left[1,2,1^{\ast},2^{\ast}\right]$, is a proper $\sigma$-class.
+We must observe that $\sigma$-set theory [5] is a theory of $\sigma$-classes, where $\sigma$-sets are characterized by axioms.
+We must also note that a proper $\sigma$-class is a $\sigma$-class that is not a $\sigma$-set.
+This difference is essential to give rise to the existence of antielements along with their respective elements.
 Definition 4.3.
-Let $A$ be an integer $\sigma$-set such that $|A|=n$, then $X$ is said to be a $\sigma$-set variable of $3^{A}$, if and only if
+Let $A$ be a integer $\sigma$-set such that $|A|=n$, then $X$ is said to be a $\sigma$-set variable of $3^{A}$, if and only if
 $X=\{x_{1},x_{2},x_{3},\ldots,x_{m}\},$
@@
 We must remember that the structure in general is not associative, therefore we cannot freely use this property, so to find the solution to the equation we must develop a previous theorem.
-To develop this theorem we will assume that for every integer $\sigma$-set $A$ then the generated space $\left\langle 2^{A},2^{A^{-}}\right\rangle=3^{A}$, and also that $3^{A}$ satisfies the conjecture 3.2.
+To develop this theorem we will assume that for every integer $\sigma$-set $A$ the generated space is $\left\langle 2^{A},2^{A^{-}}\right\rangle=3^{A}$, and also that $3^{A}$ satisfies conjecture 3.2.
 Theorem 4.7.
@@
 Proof.
-Let $A$ be an entire $\sigma$-set, $X$ be a $\sigma$-set variable of $3^{A}$ and $M\in 3^{A}$, then by lemma 4.5 we have that $X\oplus M=X\cup M$, with $X\cap M=\emptyset$.
+Let $A$ be an integer $\sigma$-set, $X$ be a $\sigma$-set variable of $3^{A}$ and $M\in 3^{A}$, then by lemma 4.5 we have that $X\oplus M=X\cup M$, with $X\cap M=\emptyset$.
 Therefore we have that
@@
 $(X\oplus M)\oplus M^{-}=X.$
-Now that theorem 4.7 has been proved, we can solve some $\sigma$-set equation for the integer $\sigma$-set $A=\{1,2\}$, since the generated space is effectively equal to $3^{A}$, that is, $\left\langle 2^{A},2^{A^{-}}\right\rangle=3^{A}$, and also $3^{A}$ is a non-associative abelian loop.
-Let $A=\{1,2\}$ be an integer set and $M,N\in 3^{A}$, then the equation
+Now, after theorem 4.7 has been proved, we can solve some $\sigma$-set equation for the integer $\sigma$-set $A=\{1,2\}$, since the generated space is effectively equal to $3^{A}$, that is, $\left\langle 2^{A},2^{A^{-}}\right\rangle=3^{A}$, and also $3^{A}$ is a non-associative abelian loop.
+Let $A=\{1,2\}$ be an integer set and $M,N\in 3^{A}$, with $M\widehat{\cap}N=\emptyset$, then the equation
 $X\oplus M=N,$
@@
 Example 4.8.
-Let $A=\{1,2\}$ be an integer $\sigma$-set, $M=\{1,2^{\ast}\}$ and $N=\{1\}$, then the equation of a $\sigma$-set variable
+Let $A=\{1,2\}$ be an integer $\sigma$-set, $M=\{1,2^{\ast}\}$ and $N=\{1\}$, with $M\widehat{\cap}N=\emptyset$, then the equation of a $\sigma$-set variable
 $X\oplus\{1,2^{\ast}\}=\{1\}$
@@
 $S=\{S_{1},S_{2}\}=\{\{2\},\{1,2\}\}.$
-Before we conclude our study of $\sigma$-sets and $\sigma$-antisets we are going to present a conjecture regarding the possible solutions of a particular type of equations of a $\sigma$-set variable.
-Conjecture 4.9.
-Let $A$ be an integer $\sigma$-set, $X$ a $\sigma$-set variable of $3^{A}$, and $M,N\in 3^{A}$, then a solution of the equation
-$X\oplus M=N,$
-is $S=N\oplus M^{-}$.
-Let us look into the following case, where $M=\{1^{*}\}$ and $N=\{1\}$, so for the equation
-$X\oplus\{1^{*}\}=\{1\},$
-there is no element $x$ such that $\{x,1^{*}\}=\{1\}$.
-This lead us into a more general case where $N\subseteq M^{-}$ or $N\supset M^{-}$, being the conjecture false, which take us into establishing the following condition in order for this type of equation being solvable.
+Note that if $M\widehat{\cap}N=\emptyset$ then the $\sigma$-set equation has a solution, but otherwise the $\sigma$-set equation has an empty solution.
+Example 4.9.
+Let $A=\{1,2\}$ be an integer $\sigma$-set, $M=\{1^{\ast}\}$ and $N=\{1\}$, with $M\widehat{\cap}N=\{1^{*}\}$, then the equation of a $\sigma$-set variable
+$X\oplus\{1^{\ast}\}=\{1\}$
+There is no solution.
+$X\oplus\{1^{\ast}\}=\{1\}\setminus\oplus\{1\},$
+$(X\oplus\{1^{\ast}\})\oplus\{1\}=\{1\}\oplus\{1\},$
+$X=\{1\},$
+which is a contradiction, because
+$\{1\}\oplus\{1^{\ast}\}=\{1\},$
+=$\emptyset=\{1\}.$
 Definition 4.10.
 A $\sigma$-set equation $X\oplus M=N$ is said to be fusionable if $M\widehat{\cap}N=\emptyset$.
-With this, let us conclude with a bounded theorem to find some solutions of the $\sigma$-set equation.
+With this in mind, let us conclude with a bounded theorem to find some solutions of the $\sigma$-set equation.
 Theorem 4.11.
@@
 Now, by using Theorem 4.11, let us first obtain
-$R$ $=$ $(M\oplus N^{-})^{-}$
+$R^{-}$ $=$ $(M\oplus N^{-})^{-}$
 $=$ ($(\{1,2,3^{*},4^{*},5,6^{*}\}\oplus\{1,2\}^{-})^{-}$
@@
 Hence $S=\{\{1,2,3,4,5^{*},6\},\{3,4,5^{*},6\}\}$ is a solution set for the fusionable equation $X\oplus M=N$.
+5 Conclusions
+One of the first conclusions we can draw is that the fusion operator $\oplus$ for $\sigma$sets is equivalent to the union operator for sets within the context of the set of parts $2^{A}$, which allows us to deduce that the fusion of $\sigma$-sets is an extension of the union for the generated space.
+The fact that the integer space $3^{A}$ presents a cardinal of power 3, is very important for the development of the theory of transfinite numbers, since in general the power set $2^{A}$ that goes to the power of 2 is used; in this way our results can serve as an impetus for the development of the theory of transfinite numbers.
+We can also conclude that the algebraic structure of the integer space $3^{1,2}$ is a loop, which leads us to conjecture that the integer space in general has a loop structure.
+This fact is relevant to $\sigma$-set theory since, if it were so, it would show that the fusion operator $\oplus$ is not associative which is relevant for solving set equations.
+As a final conclusion, we can state that we can generate $\sigma$-set equations given the existence of inverses for the fusion operator $\oplus$ in the integer space, but in the general case, solutions are not given, so a condition must be imposed on the $\sigma$-sets of the equation.
+We have not yet conducted a detailed study on the number of solutions to each set equation, leaving this study for future research.
+To see more works in which antisets or $\sigma$-antiset are used or in which equation $A\cup B=\emptyset$ is described, visit the references [1], [2], [4], [5], [6].